find quadratic equation whose roots are -3 and 5?









Answer by Edwin McCravy(20054)


It's the reverse of solving.  You start with the answers and get the
equation that has those answers.

   x = -3;    x = 5
 x-3 = 0 ;  x-5 = 0
   (x-3)(x-5) = 0
     x²-8x+15 = 0
 
Edwin